What number of molecules of CO2 will be produced if 3x10^24 molecules of O2(g) reacts?

1
Expert's answer
2022-05-02T17:04:03-0400

1 mol O2 molecules = 6.022×1023 molecules O2

Use dimensional analysis to convert molecules O2 to moles O2

3×1024 molecules O2 × (1 mol O2 molecules/6.022×1023 molecules O2) = 4.98 mol CO2


C + O2 = CO2

n(CO2) = n(O2) = 4.98 moles

N(CO2) = 4.98 • 6.022×1023 = 3×1024 molecules